(d^2-5d+6)y =x cos(2x)

1
Expert's answer
2022-02-22T08:08:47-0500

Solution:

For complementary solution:

"D^2-5D+6=0
\\\Rightarrow D^2-3D-2D+6=0
\\\Rightarrow (D-3)(D-2)=0
\\\Rightarrow D=3,D=2"

So, "y_c=c_1e^{2x}+c_2e^{3x}"

Next, we determine particular solution by method of undetermined coefficients.

The particular solution to "\frac{d^{2} y(x)}{d x^{2}}-5 \frac{d y(x)}{d x}+6 y(x)=x \cos (2 x)"  is of the form: "y_{p}(x)=a_{1} \cos (2 x)+a_{2} x \cos (2 x)+a_{3} \sin (2 x)+a_{4} x \sin (2 x)"

Solve for the unknown constants "a_{1}, a_{2}, a_{3} ,\ and\ a_{4}"  :

Compute "\frac{d y_{p}(x)}{d x} :

 \frac{d y_{p}(x)}{d x}=\frac{d}{d x}\left(a_{1} \cos (2 x)+a_{2} x \cos (2 x)+a_{3} \sin (2 x)+a_{4} x \sin (2 x)\right)"

"\begin{aligned}

=&-2 a_{1} \sin (2 x)+a_{2} \cos (2 x)-\\

& 2 a_{2} x \sin (2 x)+2 a_{3} \cos (2 x)+2 a_{4} x \cos (2 x)+a_{4} \sin (2 x)

\end{aligned}"

Compute "\frac{d^{2} y_{p}(x)}{d x^{2}} :"

"\begin{aligned}

\frac{d^{2} y_{p}(x)}{d x^{2}}=& \frac{d^{2}}{d x^{2}}\left(a_{1} \cos (2 x)+a_{2} x \cos (2 x)+a_{3} \sin (2 x)+a_{4} x \sin (2 x)\right) \\

=&-4 a_{1} \cos (2 x)-4 a_{2} x \cos (2 x)-\\

& 4 a_{2} \sin (2 x)-4 a_{3} \sin (2 x)+4 a_{4} \cos (2 x)-4 a_{4} x \sin (2 x)

\end{aligned}"

Substitute the particular solution "y_{p}(x)"  into the differential equation:

"\begin{aligned}

&\frac{d^{2} y_{p}(x)}{d x^{2}}-5 \frac{d y_{p}(x)}{d x}+6 y_{p}(x)=x \cos (2 x) \\

&-4 a_{1} \cos (2 x)-4 a_{2} x \cos (2 x)-4 a_{2} \sin (2 x)-4 a_{3} \sin (2 x)+ \\

&4 a_{4} \cos (2 x)-4 a_{4} x \sin (2 x)-5\left(-2 a_{1} \sin (2 x)+a_{2} \cos (2 x)-\right. \\

&\left.2 a_{2} x \sin (2 x)+2 a_{3} \cos (2 x)+2 a_{4} x \cos (2 x)+a_{4} \sin (2 x)\right)+ \\

&6\left(a_{1} \cos (2 x)+a_{2} x \cos (2 x)+a_{3} \sin (2 x)+a_{4} x \sin (2 x)\right)=x \cos (2 x)

\end{aligned}"

Simplify:

"\begin{aligned}

&\left(2 a_{1}-5 a_{2}-10 a_{3}+4 a_{4}\right) \cos (2 x)+\left(2 a_{2}-10 a_{4}\right) x \cos (2 x)+ \\

&\quad\left(10 a_{1}-4 a_{2}+2 a_{3}-5 a_{4}\right) \sin (2 x)+\left(10 a_{2}+2 a_{4}\right) x \sin (2 x)=x \cos (2 x)

\end{aligned}"

Equate the coefficients of "\cos (2 x)"  on both sides of the equation:

"2 a_{1}-5 a_{2}-10 a_{3}+4 a_{4}=0"

Equate the coefficients of "x \cos (2 x)"  on both sides of the equation:

"2 a_{2}-10 a_{4}=1"

Equate the coefficients of "\sin (2 x)"  on both sides of the equation:

"10 a_{1}-4 a_{2}+2 a_{3}-5 a_{4}=0"

Equate the coefficients of "x \sin (2 x)"  on both sides of the equation:

"10 a_{2}+2 a_{4}=0"

Solve the system:

"\begin{aligned}

&a_{1}=-\frac{5}{169} \\

&a_{2}=\frac{1}{52} \\

&a_{3}=-\frac{73}{1352} \\

&a_{4}=-\frac{5}{52}

\end{aligned}"

Substitute"a_{1}, a_{2}, a_{3} ,\ and\ a_{4}\ into\ y_{p}(x)=\cos (2 x) a_{1}+x \cos (2 x) a_{2}+\sin (2 x) a_{3}+ x \sin (2 x) a_{4}"  :

"y_{p}(x)=-\frac{5}{169} \cos (2 x)+\frac{1}{52} x \cos (2 x)-\frac{73 \sin (2 x)}{1352}-\frac{5}{52} x \sin (2 x)"

The general solution is:

Answer:

"\begin{aligned}

&y(x)=y_{\mathrm{c}}(x)+y_{p}(x)=-\frac{5}{169} \cos (2 x)+ \\

&\frac{1}{52} x \cos (2 x)-\frac{73 \sin (2 x)}{1352}-\frac{5}{52} x \sin (2 x)+c_{1} e^{2 x}+c_{2} e^{3 x}

\end{aligned}"